# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

### v2

@@ III-E Summary of the proposed controller design
The proposed data-driven design procedure can be summarized as follows. In Procedure 1, $C_{FO}\left(s;\theta\right)$ is an FO controller with the tunable parameter $\theta$. Procedure 1 can handle various types of controllers and reference models; the present approach is highly versatile and flexible.
@@
    Set $C\left(z;\theta_{0}\right)$ , where $\theta_{0}$ is the initial guess of the controller parameter.
@@
    Acquire the initial data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ through a closed-loop experiment using $C\left(z;\theta_{0}\right)$ and $r_{[0:N]}^{D}$.
@@
    Solve the optimization problem (5). The solution of (5) is denoted as $\theta^{\ast}$.
  7. 7.
@@
In practice, the IO approximation and discretization have to be carried out so as to implement an FO controller [17, 55]. The IO approximation and discretization may affect the closed-loop performance and stability [56]. Unlike conventional approaches, the proposed design scheme explicitly handles the IO-approximated and discretized controller. That is, the proposed approach evaluates the resultant closed-loop performance and stability for the IO-approximated and discretized, i.e., ready-to-implement, controller, showing the reliability of the proposed approach. $\blacksquare$
IV Numerical examples
[…]
@@ IV Numerical examples
where $K_{p},K_{i},K_{d},K_{fp},K_{fi},K_{fd},\lambda,\mu,\tau_{i},\tau_{f}\in\mathbb{R}_{+}\cup\left\{0\right\}$. The FOPID controller is a typical example of the FC-based control scheme [61]. In this study, we set $\tau_{i}=\tau_{f}=t_{s}$. The tuning conditions are listed in Table I. The IO approximation is performed via the Oustaloup recursive filter (order: 5; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-4},10^{4}\right)\;\mathrm{rad/s}$) for both $C_{FOPID}\left(z;\theta\right)$ and $M_{ref}\left(z\right)$. In the initial data correction, the simulation time is $40\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data for the controller tuning are the same to each other in the FO-PID controller and the IO-PID controller, since $C_{IOPID}\left(z;\theta_{0}\right)=C_{FOPID}\left(z;\theta_{0}\right)$. Fig. 4 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ for tuning both $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$.
Table II summarizes the tuning results for $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$ provided by the proposed approach (iso-IDFRIT). Table III lists the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by the tuned controllers. Fig. 5 demonstrates the control results obtained by the resultant controllers. The Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ are shown in Fig. 6. In Fig. 5, the black line indicates the setpoint reference input. In Figs. 5 and 6, the green, blue, and red lines demonstrate the results due to the reference model ($M_{ref}\left(z\right)$ in Fig. 5 and $L_{flat}\left(z\right)$ in Fig. 6), $C_{IOPID}\left(z;\theta^{\ast}\right)$, and $C_{FOPID}\left(z;\theta^{\ast}\right)$, respectively. Tables II and III and Figs. 5 and 6 clearly show that the proposed approach successfully provides the IO- and FO-PID controllers yielding the desired closed-loop property. In particular, the closed-loop system due to the FO-PID controller achieves the closer specifications ($\phi_{m}$, $\omega_{c}$, the flat-phase) to the desired ones than that due to the IO-PID controller.
TABLE I: Conditions for Example 1
Search range $K_{p},K_{i},K_{d}\in\begin{bmatrix}0&5\\ \end{bmatrix}$
$J\left(\theta_{0}\right)$ $9.8820\times 10^{2}$
TABLE II: Tuning results (Example 1)
@@
TABLE III: Frequency domain property provided by the tuned controllers (Example 1).
@@
Figure 5: Control results due to the tuned controllers (Example 1). The green and red lines almost overlap
Figure 6: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ (Example 1): (a) overview, and (b) enlarged phase plots around $1\,\mathrm{rad/s}$.
@@
IV-B Example 2: Soft robot (superiority over the conventional model-based approach)
One representative virtue of the proposed approach is the freedom from the necessity of the plant model and the model reduction. This feature allows the proposed approach to achieve better control performance and robustness than the reduced-model-based controller design scheme. Note that almost all conventional analytical and graphical design methodologies rely on the reduced order model, as discussed in Section I-B. We demonstrate the superiority and robustness of the proposed approach, compared to the traditional iso-damping controller design based on the reduced model. Throughout this example, the Oustaloup recursive filter is employed for the IO approximation (order: 7; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-3},10^{6}\right)\;\mathrm{rad/s}$) for both controllers and reference models. We discretize the plant, the reference model, and the controller in the sampling time $t_{s}=10^{-2}\,s$.
@@ IV-B Example 2: Soft robot (superiority over the conventional model-based approach)
where $K_{fp},K_{fi},\lambda\in\mathbb{R}_{+}\cup\left\{0\right\}$. The tuning condition is listed in Table IV. Fig. 8 describes the initial input/output data obtained using $C_{FOPI}\left(z;\theta_{0}\right)$, where the simulation time for initial data collection is $2\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data is collected on the basis of the actual plant $P\left(z\right)$ in (21).
@@
[…]
TABLE IV: Conditions for Example 2.
$J\left(\theta_{0}\right)$ $3.8646\times 10^{1}$
TABLE V: Tuning result of the proposed approach (Example 2).
@@
TABLE VI: Frequency domain property with the actual plant $P\left(z\right)$ (Example 2).
@@
Figure 10: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{proposed}\left(z;\theta^{\ast}\right)$ (Example 2): (a) overview, and (b) enlarged phase plots around $12\,\mathrm{rad/s}$.
@@ V Discussion
Unlike traditional analytical and graphical approaches, the proposed approach does not require plant model experiments. Neither repeated experiments nor frequency response computations is necessary, whereas such procedures are indispensable for conventional data-driven and optimization-based approaches. In the present approach, the desired controller is automatically designed via solving the optimization problem defined based on only one-shot input/output data; the BIBO stability from a bounded reference input to a bounded output is explicitly considered for the IO-approximated and discretized (i.e., ready-to-implement) controllers. Hence, the proposed approach is a simple, practical, and reliable controller design technique so as to achieve the iso-damping property.
@@
In this study, we assume that the reference model is given by the designer. That is, the designer must give the appropriate reference model. An important future direction may be the simplification of the design procedure for the reference model. For example, the flat-phase property is frequency-domain specification, whereas it is more straightforward and intuitive to design the reference model in the time-domain. The Auto-DDC approach [62] may be effective for addressing this issue. One study has proposed the data-driven tuning of parameters of both the controller and the reference model [63]. In the future, we will examine the Auto-DDC strategy and the simultaneous tuning technique of the controller and the reference model in iso-IDFRIT.

### v3

@@ III-E Summary of the proposed controller design
Procedure 1 summarizes the proposed data-driven design method. In Procedure 1, $C_{FO}\left(s;\theta\right)$ is an FO controller with the tunable parameter $\theta$. Procedure 1 can handle various types of controllers and reference models; the present approach is highly versatile and flexible. Figure 4 describes the solution procedure of the optimization problem for the proposed controller design approach. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)$, which is used to compute $J\left(\theta\right)$, can be generated straightforwardly by filtering $u_{[0:N]}^{D}$ by $\left(C\left(z;\theta\right)\right)^{-1}$ and adding it to $y_{[0:N]}^{D}$, as shown in (4).
@@
    Acquire the initial data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ through an open-loop or closed-loop experiment. If these data are corrupted by noise, mitigate the data noise using a denoising technique (e.g., low-pass filtering, the total variation denoising [54]).
@@
    Solve the optimization problem (10) for the reference input $r_{[0:N]}^{D}$. The solution of (10) is denoted as $\theta^{\ast}$.
@@
We can interpret the loss function $J\left(\theta\right)$ from the viewpoint of the frequency domain. Specifically, since $J\left(\theta\right)=\sum_{n=0}^{N}\left\lvert\left(t_{k}\left(\theta\right)-m_{k}^{ref}\right)\ast r_{k}^{D}\right\rvert^{2}$,
where $\Phi_{rD}\left(e^{j\omega}\right)$ is the spectral density of the reference input $r_{[0:\infty]}^{D}$. Therefore, the minimization of $J\left(\theta\right)$ corresponds to the minimization of the mismatch of the frequency responses $T\left(e^{j\omega};\theta\right)$ and $M_{ref}\left(e^{j\omega}\right)$ with the frequency weighting function $\Phi_{rD}\left(e^{j\omega}\right)$.
Remark 8.
In practice, the IO approximation and discretization have to be carried out in order to implement an FO controller [17, 59]. The IO approximation and discretization may affect the closed-loop performance and stability [60]. Unlike conventional approaches, the proposed design scheme explicitly handles the IO-approximated and discretized controller. Specifically, the proposed approach evaluates the resultant closed-loop performance and stability for the IO-approximated and discretized, i.e., ready-to-implement, controller, showing the reliability of the proposed approach.
Figure 4: Solution procedure of the optimization problem for the proposed controller design approach. The loss function $J\left(\theta\right)$ shown in (LABEL:Eq_Loss_iso-IDFRIT) can be minimized without iteratively conducting closed-loop control tests.
IV Validation
The effectiveness of the proposed approach is demonstrated in this section. Throughout this study, the Oustaloup recursive filter [61, 62] is employed for the IO approximation $\mathrm{f2i}$ for both the reference model and the controller; we use the Tustin method for all $\mathrm{c2d}$ schemes. The design and IO approximation of the FO transfer function are performed via the FOMCON toolbox [63]. The iso-damping reference model is designed using the BITF shown in (3), where the integral order and the gain crossover frequency are denoted as $\gamma$ and $\omega_{c}$ $\mathrm{[rad/s]}$, respectively. Here, $\gamma=2\left(1-\frac{\phi_{m}}{\pi}\right)$, where $\phi_{m}$ $\mathrm{[rad/s]}$ represents the phase margin [20]. That is, $M_{ref}\left(z\right)=L_{flat}\left(z\right)\left\{1+L_{flat}\left(z\right)\right\}^{-1}$, where $L_{flat}\left(z\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(L_{Bode}\left(s;\psi\right)\right)$ for $\psi=\begin{bmatrix}\phi_{m}&\omega_{c}\\ \end{bmatrix}$. We use the particle swarm optimization to solve (10), which is implemented in the MATLAB Global Optimization Toolbox (particleswarm). Note that the plant model is not used at all for the proposed controller design approach in all examples.
Remark 9.
The PSO algorithm is used in this study because the nature of the PSO algorithm is well-suited to the optimization problem (10). In particular, the problem (10) is non-convex, which requires a careful search over the solution space. Thus, population-based optimization strategies, including PSO, are well-suited for addressing (10). Notably, we can easily use population-based algorithms as a solver for (10), as the value of $J\left(\theta\right)$ can be evaluated without conducting closed-loop experiments with the controllers under test. This feature greatly reduces the cost to evaluate the objective function value compared with conventional optimization-based approaches that must evaluate it by actually implementing each controller to be tested into the real-world system and conducting closed-loop control experiments. Here, PSO is such a common population-based algorithm that it has already been implemented in commercial numerical software. Hence, using PSO as the solver in the verification examples shows the high practicality of the proposed approach.
@@ IV Validation
TABLE II: Conditions for Example 1
Search range $K_{p},K_{i},K_{d}\in\begin{bmatrix}0&5\\ \end{bmatrix}$
$J\left(\theta_{0}\right)$ $9.8820\times 10^{2}$
@@
where $K_{p},K_{i},K_{d},K_{fp},K_{fi},K_{fd},\lambda,\mu,\tau_{i},\tau_{f}\in\mathbb{R}_{+}\cup\left\{0\right\}$. The FO-PID controller is a typical example of the FC-based control scheme [65]. In this study, we set $\tau_{i}=\tau_{f}=t_{s}$.
The tuning conditions are listed in Table II. The IO approximation is performed via the Oustaloup recursive filter (order: 5; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-4},10^{4}\right)\;\mathrm{rad/s}$) for both $C_{FOPID}\left(z;\theta\right)$ and $M_{ref}\left(z\right)$. Fig. 5 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ for tuning both $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$. The input/output data are collected through the closed-loop control simulation with the unit step as the reference input $r_{[0:N]}^{D}$ and the initial controller parameter $\theta_{0}$ shown in Table II, where the simulation time is $40\,\mathrm{s}$. Note that the initial data for the controller tuning are the same to each other in the FO-PID controller and the IO-PID controller, since $C_{IOPID}\left(z;\theta_{0}\right)=C_{FOPID}\left(z;\theta_{0}\right)$.
Table III summarizes the tuning results for $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$ provided by the proposed approach (i.e., Iso-IDFRIT). Table IV lists the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by the tuned controllers. Fig. 6 demonstrates the control results obtained by the resultant controllers. The Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ are shown in Fig. 7. In Fig. 6, the black line indicates the setpoint reference input. In Figs. 6 and 7, the green, blue, and red lines demonstrate the results due to the reference model ($M_{ref}\left(z\right)$ in Fig. 6 and $L_{flat}\left(z\right)$ in Fig. 7), $C_{IOPID}\left(z;\theta^{\ast}\right)$, and $C_{FOPID}\left(z;\theta^{\ast}\right)$, respectively. Tables III and IV and Figs. 6 and 7 clearly show that the proposed approach successfully provides the IO- and FO-PID controllers yielding the desired closed-loop property. Note that the closed-loop system due to the FO-PID controller achieves the closer specifications ($\phi_{m}$, $\omega_{c}$, the flat-phase) to the desired ones than that due to the IO-PID controller.
TABLE III: Tuning results (Example 1)
@@
TABLE IV: Frequency domain property provided by the tuned controllers (Example 1).
@@
Figure 6: Control results due to the tuned controllers (Example 1). The green and red lines almost overlap.
Figure 7: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ (Example 1): (a) overview, and (b) enlarged phase plots around $1\,\mathrm{rad/s}$.
@@
IV-B Example 2: Soft robot (superiority over the conventional approaches)
One representative virtue of the proposed approach is the freedom from the necessity of the plant model and the model reduction. This feature allows the proposed approach to achieve better control performance and robustness than the reduced-model-based controller design scheme. Note that almost all conventional analytical and graphical design methodologies rely on the reduced order model, as discussed in Section I-B. We demonstrate the superiority and robustness of the proposed approach, compared to the traditional iso-damping controller design based on the reduced model. Throughout this example, the Oustaloup recursive filter is employed for the IO approximation (order: 7; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-3},10^{6}\right)\;\mathrm{rad/s}$) for both controllers and reference models. We discretize the plant, the reference model, and the controller in the sampling time $t_{s}=10^{-2}\,\mathrm{s}$.
@@ IV-B Example 2: Soft robot (superiority over the conventional approaches)
where $K_{fp},K_{fi},\lambda\in\mathbb{R}_{+}\cup\left\{0\right\}$. The tuning condition is listed in Table V. Fig. 9 shows the initial input/output data used for the proposed approach. These were collected through the closed-loop control simulation with $C_{FOPI}\left(z;\theta_{0}\right)$ as the initial controller and the unit step as the reference input $r_{[0:N]}^{D}$, where the simulation time was $2\,\mathrm{s}$. Note that the initial data is collected on the basis of the actual plant $P\left(z\right)$ in (27).
TABLE V: Conditions for Example 2.
$J\left(\theta_{0}\right)$ $3.8646\times 10^{1}$
@@
TABLE VI: Tuning result of the proposed approach (Example 2).
@@
TABLE VII: Frequency domain property with the actual plant $P\left(z\right)$ (Example 2).
@@
Figure 11: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{proposed}\left(z;\theta^{\ast}\right)$ (Example 2): (a) overview, and (b) enlarged phase plots around $12\,\mathrm{rad/s}$.
The tuning result due to the proposed approach is summarized in Table VI. Table VII compares the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by $C_{iso-m}\left(z\right)$ and $C_{proposed}\left(z\right)$. Fig. 10 demonstrates the control results provided by $C_{iso-m}\left(z\right)$ and $C_{proposed}\left(z\right)$; Fig. 11 shows the Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z\right)$, and $P\left(z\right)C_{proposed}\left(z\right)$. The setpoint reference is indicated by the black line in Fig. 10. In Figs. 10 and 11, the results due to the reference model, $C_{iso-m}\left(z\right)$, and $C_{proposed}\left(z\right)$ are described by the green, cyan, and red lines, respectively. Note that the results in Table VII and Figs. 10 and 11 are computed on the basis of not the reduced order plant $\mathrm{c2d}\left(P_{reduced}\left(s\right)\right)$ but the actual controlled plant $P\left(z\right)$. Results in Table VII and Figs. 10 and 11 clearly demonstrate that $C_{proposed}\left(z\right)$ achieves the closed-loop properties ($\phi_{m}$, $\omega_{c}$, the flat-phase) much closer to the desired ones than $C_{iso-m}\left(z\right)$.
@@
To further highlight its advantage, the proposed approach is compared with IFT as a traditional data-driven FO controller tuning technique. Numerous studies have reported the effectiveness of the IFT approach for tuning FO controllers, e.g., [46]. In this example, IFT seeks $\theta=\begin{bmatrix}K_{fp}&K_{fi}&\lambda\\ \end{bmatrix}$ that minimizes $E\left(\theta\right)=\left\lVert T\left(z;\theta^{\ast}\right)r_{[0:N]}^{D}-M_{ref}\left(z\right)r_{[0:N]}^{D}\right\rVert_{2}^{2}$, which is the same MR control problem as the proposed approach. Specifically, the IFT approach finds the minimizer of $E\left(\theta\right)$ by iteratively conducting closed-loop tests and updating $\theta$ via PSO; PSO is used to ensure fair comparison with the proposed approach. As a result, the IFT method provided $\theta=\begin{bmatrix}8.8086\times 10^{-1}&3.8808&4.7498\times 10^{-1}\\ \end{bmatrix}$ , which is the same as that obtained by the proposed approach. This result stems from the fact that the proposed objective function (LABEL:Eq_Loss_iso-IDFRIT) is an exact reformulation of $E\left(\theta\right)$ based on a single set of input and output data, as proved in Theorem 1. This fact indicates that the proposed approach has a significant advantage over IFT for tuning FO controllers in that the required number of the closed-loop tests is reduced.
IV-C Example 3: Experimental verification
[…]
Figure 13: Experimental setup of Example 3. The control input $u$ is the command value of the voltage applied to drive the motor. The output $y$ is the voltage generated by the generator, which is proportional to the rotational speed of the generator.
In this example, data for the proposed approach is collected through an open-loop control experiment using the proportional controller $C_{ini}=2.5$, as shown in Fig. 14. The data collection begins when a step reference is applied to the controlled plant in a steady-state i.e., the motor is rotating steadily. As the output data is corrupted by measurement noise in real-world systems, the output data is mitigated using the $L_{2}$ total variation denoising technique [54]:
$\mathrm{vec}\left(y_{[0:N]}^{D}\right)=\left(I+\nu\chi^{\top}\chi\right)^{-1}\mathrm{vec}\left(\hat{y}_{[0:N]}^{D}\right)$ (31)
$\chi\coloneqq\begin{bmatrix}-1&1&0&\ldots&0\\ 0&-1&1&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\ldots&0&-1&1\\ 0&\ldots&0&0&0\\ \end{bmatrix}$ (32)
where $\nu\in\mathbb{R}_{+}$ (we set $\nu$ to $2$ in this example), $I$ is the identity matrix with the compatible dimension, $\hat{y}_{[0:N]}^{D}$ is the noisy output data shown in Fig. 14, and $y_{[0:N]}^{D}$ is the denoised output data. The $L_{2}$ total variation denoising technique is often employed to handle noisy data in the data-driven controller design framework, owing to its computational efficiency. The details of this denoising approach have been discussed in [54]. Figure 15 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ and the setpoint reference $r_{[0:N]}^{D}$ for the proposed approach.
Figure 14: Sketch of the initial data collection procedure. The data is collected through an open-loop control experiment using the proportional controller $C_{ini}=2.5$. The data collection begins when a step reference is applied to the controlled plant in a steady-state (i.e., the motor is rotating steadily).
The tuning result provided by the proposed approach is $\theta^{\ast}=\begin{bmatrix}K_{fp}&K_{fi}&\lambda\\ \end{bmatrix}=\begin{bmatrix}1.2026&4.9931&1.0911\\ \end{bmatrix}$. Figure 16 shows the control results obtained using $C_{FOPI}\left(z;\theta^{\ast}\right)$, i.e., the FO-PI controller tuned by the proposed approach. The black dashed-dotted line represents the reference input, which specifies the set point. The red, green, and magenta lines show the control results for $P\left(z\right)$ (nominal plant), $1.5P\left(z\right)$ (plant with a $50\%$ increase in gain), and $0.5P\left(z\right)$ (plant with a $50\%$ decrease in gain), respectively. The blue line represents the response of the reference model $M_{ref}\left(z\right)$. The comparison of the blue and red lines demonstrates that the proposed approach successfully yields the FO-PI controller that achieves the closed-loop response similar to the desired response given by $M_{ref}\left(z\right)$. As shown by the comparison of the red, green, and magenta lines, this tuning result achieves the iso-damping property. Thus, we validate the effectiveness of our proposed controller design approach for real-world control systems.
Figure 15: Input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ and setpoint reference $r_{[0:N]}^{D}$ for proposed controller tuning. The output data is denoised using the $L_{2}$ total variation denoising.
Figure 16: Control results due to the FO-PI controller tuned by the proposed approach. The comparison of the red, green, and magenta lines verifies the robustness of the overshoot amount against variations in plant gain.
@@ V Discussion
Unlike traditional analytical and graphical approaches, the proposed approach does not require plant models. Neither repeated experiments nor frequency response computations is necessary, whereas such procedures are indispensable for conventional data-driven and optimization-based approaches. In the present approach, the desired controller is automatically designed via solving the optimization problem defined based on only one-shot input/output data; the BIBO stability from a bounded reference input to a bounded output is explicitly considered for the IO-approximated and discretized (i.e., ready-to-implement) controllers. Hence, the proposed approach is a simple, practical, and reliable controller design technique to achieve the iso-damping property.
@@
Example 3 experimentally verifies the effectiveness of the proposed controller design approach in real-world systems. Real-world systems, including the one in this example, pose various challenges. For example, noise corrupting the measurement data may affect the data-based controller design approaches. In practice, the proposed approach is applicable to real-world systems by mitigating the noise in the data for controller design using well-established denoising techniques. In our experimental verification, the $L_{2}$ total variation denoising [54] is adopted to mitigate the measurement noise. Note that this denoising technique is so simple and easy-to-use that incorporating this denoising procedure does not unduly complicate the proposed approach. Regarding issues related to the implementation of the proposed approach, mathematically proving BIBO stability is difficult in practical real-world systems due to unmodeled uncertainties and unexpected noise. Thus, in practice, it is important to reasonably infer the closed-loop properties before implementing the controller. From this perspective, the proposed approach yields a reliable controller, as we can reasonably infer that a controller parameter $\theta$ that minimizes $J\left(\theta\right)$ will yield $T\left(z;\theta\right)$ such that it is BIBO stable (see Remark 5). In fact, in Example 3, the FOPI controller tuned by the proposed approach achieves good control performance and robustness without causing instability. Consequently, the proposed approach is a simple and effective technique for designing an iso-damping robust controller for real-world control systems. We will mathematically evaluate the effects of various uncertainties and constraints in real-world systems on the proposed approach in the future.
In this study, we assume that the reference model is given by the designer. That is, the designer must give the appropriate reference model. An important future direction may be the simplification of the design procedure for the reference model. For example, the flat-phase property is a frequency-domain specification, whereas it is more straightforward and intuitive to design the reference model in the time-domain. The Auto-DDC approach [66] may be effective for addressing this issue. One study has proposed the data-driven tuning of parameters of both the controller and the reference model [67]. In the future, we will examine the Auto-DDC strategy and the simultaneous tuning technique of the controller and the reference model in iso-IDFRIT.